OpenStudy (anonymous):

Franklin's hair grew one-sixth inches in one week. If his hair continues to grow at this rate, about how many inches would his hair grow in one year, or fifty-two weeks? 1 3 6 9

OpenStudy (anonymous):

@Innerlines

OpenStudy (anonymous):

@haiimawesome

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

1/6 = x/52 6x=52

OpenStudy (anonymous):

so divide 52 by 6

OpenStudy (anonymous):

It seems that for this question you have to multiply 1/6 by 52 or create a type of proportion such as 1/6=x/52.

OpenStudy (anonymous):

somwehere aroun eight

OpenStudy (anonymous):

if it is a decimal, round

OpenStudy (anonymous):

8.like 16

OpenStudy (anonymous):

that should be 8.6666666667 so round up